Graph a line with a slope of -5 that contains the point (-3,-4)

Found 2 solutions by mananth, josgarithmetic:
Answer by mananth(16946)

Graph a line with a slope of -5 that contains the point (-3,-4)
Let's find theY intercept
y=mx+b is the equation in slope intercept form
plug x,y,m
-4= (-5)(-3)+ b
-4 = 15+b
b=-19
The line intercepts the Y axis at (0,-19)
The equation of the line will be y =(-5)x-19
y=-5x-19
you can use a graphing calculator
OR
plot points (-3,-4) and (0,-19) and draw the required line







Answer by josgarithmetic(39617)

You have the simply-given information to graph the line. What to do is more than clear enough.